OpenStudy (anonymous):

fine the probability of dealing a five-card hand that contains the ace, king, queen, jack and ten of the same suit from a deck of 52 cards

OpenStudy (whpalmer4):

Pretty darn small :-)

OpenStudy (agent0smith):

There's four ways this can happen: all hearts/spades/clubs/diamonds Total number of possible 5 card hands is (from 52 cards, choose 5): 52C5 = 2598960 There are only four possible desirable outcomes: AKQJT hearts, AKQJT diamonds, AKQJT spades, AKQJT clubs. Probability = # desired outcomes/total possible outcomes.

OpenStudy (anonymous):

the answer is its 1/649740

OpenStudy (agent0smith):

Try what I gave you and simplify the fraction.

OpenStudy (agent0smith):

\[P = \frac{ 4 }{ 2598960} = ? \]

OpenStudy (anonymous):

how u come with 4? how u get that?

OpenStudy (agent0smith):

Read what I wrote above again.

OpenStudy (anonymous):

yeah oi got it thanks!

OpenStudy (whpalmer4):

\[52C5 = \frac{52!}{(52-5)! 5!} = \frac{52*51*50*49*48 }{5*4*3*2*1} = 2598960\]

OpenStudy (agent0smith):

Now you know you odds of being dealt a royal flush (in 5 card draw) :)

OpenStudy (whpalmer4):

Lesson: if you want a royal flush, bring a few extra cards ;-)

OpenStudy (agent0smith):

haha. Well it's more likely in texas hold 'em, I think I've hit 3ish playing hold 'em.

OpenStudy (whpalmer4):

how many decks were in play?

OpenStudy (agent0smith):

One... why would you play with more than one deck?

OpenStudy (agent0smith):

I should note that I've played several hundred thousand hands of hold 'em :P

OpenStudy (agent0smith):

I just checked, and I have 3 royal flushes, and 11 more straight flushes, out of 597,000 hands.

OpenStudy (skullpatrol):

Here is an interesting way to think of probability, what is the chance that a head will turn up after two flips of a coin?

OpenStudy (agent0smith):

I'm guessing you mean what's the chance of getting a head on the 2nd flip... 50%. What's the chance of getting a head on the 10th flip, if the previous nine flips were heads?

OpenStudy (skullpatrol):

What's the total chance that a head will turn up either on the first flip or the second flip?

OpenStudy (agent0smith):

If you mean an exclusive 'or' ie not allowing HH, then 50%. 75% with an inclusive 'or': HH, HT, TH = 3 desired outcomes HH, TT, HT, TH = 4 total possible outcomes

OpenStudy (agent0smith):

I think you meant an exclusive or, that's usually the standard in maths. So 50%.

OpenStudy (skullpatrol):

I meant the inclusive or, because that gives the strange result.

OpenStudy (agent0smith):

But it's not too strange if you think of it as "if you toss two coins, what's the chance of getting at least one head". The probability of at least one head will of course increase with the number of coin tosses. eg if you tossed a coin 10 times, the chance of zero heads coming up is 0.5^10.

OpenStudy (skullpatrol):

You're right! The strangeness is gone if you think of it as tossing two coins once :)

OpenStudy (agent0smith):

Or if you think of it as very many coin tosses. If you toss a coin 100 times, you'd expect to see a few heads show up!